OpenStudy (anonymous):

What is the resulting equation when the expression for y in the second equation is substituted into the first equation? 3x + y = 1 y = 6 - 4x -x + 6 = 1 x + 6 = 1 7x + 6 = 1

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (trojanpoem):

Replace each y in the first equation with (6-4x) 3x + 6 -4x = 1 -x + 6 = 1

OpenStudy (trojanpoem):

you could also solve for x, and check it in each equation.

OpenStudy (anonymous):

okay thanks:)